answersLogoWhite

0

🎒

Probability

The probability of a certain event is a number expressing the likelihood that a specific event will occur, expressed as the ratio of the number of actual occurrences to the number of possible occurrences. In mathematics, it is a measure of how often an event will happen and is the basis of statistics.

14,643 Questions

What is the probability of getting a two on one dice?

Total number of possible rolls with 2 dice = 36.

Total number of rolls that are doubles = 6.

Probability of rolling doubles= 6/36 = 1/6 = (16 and 2/3) percent .

What is the probability of rolling a number greater than 2 and a coin landing on heads if you roll a number cube and flip a coin?

The faces greater than 2 are 3, 4, 5 and 6 ; that's 4/6 = 2/3

A flipped coin will be Heads 1/2 and Tails 1/2

Combining these we have a probability of 1/2 times 2/3 = 1/3

So the probability of the combined events is 1/3 or 33.3(recurring)%

How many i's in 52 card deck?

All the cards that are diamonds have an i in, so that's 13. Of the other suits, there are 3 fives, 3 sixes, 3 eights, 3 nines and 3 kings. So in total, there are 28 cards with an i in. The probability of getting a card with an i is therefore 28/52. This can be simplified to 7/13

What is the experimental probability of rolling one die 100 times and getting a 1 or a 2 39 times?

Experimental probability depends on each trial. It is not known until you conduct the trial. You are probably talking about theoretical probability in the question.

The solve this problem, simplify it...

The probability of rolling a 1 or a 2 in a standard six sided die is 1 in 3, or about 0.3333. Remember this, because I'm going to change the problem.

Think about throwing a coin. What is the probability of throwing 39 heads in 100 coin tosses? Its the same problem, and I'll show this at the end.

The probability of throwing a head is 1 in 2, or 0.5. The probability of throwing 39 heads in 100 coin tosses is more difficult.

The real question is, what is the number of combinations of 100 things taken 39 at a time? Sound familiar? It better. Each of those 39 events corresponds to a head, and the 100 things is the number of coin tosses. The answer is that the number of combinations of 100 things taken 39 at a time is 100! / (39! (100-39)!), or about 9.0139 x 1027.

Since probability is the number of permutations of the desired outcome divided by the number of permutations of the possible outcome, simply divide the above number by 100!, or 9.3326 x 10157, getting about 9.6585 x 10-131.

Now, back to the original problem. We were talking about a six-sided die, and wanting a sample space of 6, but coins have a a sample space of 2. Reconsider and you find that the actual sample space is 3, because we want a 1 or a 2, so multiply 9.6585 x 10-131 by 2/3 to get 6.439 x 10-131. That is the probability of rolling a 1 or a 2 exactly 39 times in 100 rolls.

From Rafaelrz:

You have two possible events in a single throw of the die: A, die comes out 1 or 2; B, die comes out 3,4,5,or 6.

Probability of events are; P(A) is 1/3, P(B) is 2/3.

If you want 39 'A' events in a trial of 100 throws, the number of ways this events

can happen ( in which number of throw they appear ) is given by the number of

combinations above calculated nCr (n is 100, r is 39) with the result of about

9.01392403 x 1027.

Then, the probability of getting 39 'A' events in a trial of 100 throws is:

100C39 x [P(A)]39 x [P(B)]61 equal to 9.01392403 x 1027 x (1/3)39 x (2/3)61

equal to 0.040329065 ( Theoretical probability )

The probability of an armed confrontation between the US and the Soviets after the Vietnam war?

With the exceptions of Korea, the Cuban missile crisis, and Vietnam, the confrontation status between the US and the USSR remained about the same thru-out the Cold War.

What are the disadvantages of sampling?

You do not have all the information and so your conclusions are based on approximations.

What is the probality of rolling a sum of 5 or 6 when rolling two cubes?

-- The first cube can show any one of 6 numbers. For each of those, the second cube can also show any one of 6 numbers. So there are (6 x 6) = 36 ways the two cubes can fall. -- 5 can be rolled in any of these ways: 1-4, 2-3, 3-2, 4-1 = 4 ways -- 6 can be rolled in any of these ways: 1-5, 2-4, 3-3, 4-2, 5-1 = 5 ways -- The roll can be either 5 or 6 in (4 + 5) = 9 ways. -- Probability = (number of successes) / (number of possibilities) = (9 / 36) = 1/4 = 25 %

Draw a single card from a 52 card deck and not an ace?

Probability not ace is 1 minus probability of an ace which is 4/52. So, 1 - 4/52 is 48/52 or 12/13.

Does the Erdos Number follow the Benford Law?

No, it does not. The Benford Law requires the numbers to span several orders of magnitude, Erdos numbers have a very small range.

What is an 0001 probability level?

The answer depends on the underlying distribution.

How many cubic centimeters in 27.8 ml?

CC's and ml are the same value; therefore 27.8 ml = 27.8 cc's.

What is a probability distribution for rolling 3 dice?

The answer depends on the probability distribution of WHAT variable. The variable could be the sum or the product of the three numbers, the maximum, minimum, the mean, median, number of 3s, number of primes, and so on.

Can the probability of paternity be 0 percent?

If you have a sperm amount less than 3 ml the probability decreases.

No, sorry to say but all it takes is one sperm. On the other hand when testing for paternity the test is never 0, it can be close to it but never zero, due to the + or - error in the mathematical formula used.

How is probability used in business management?

Probability is often used to develop projections, forecasts and budgets.

What was the longest hole-in-one?

The Longest Hole-In-One ever recorded is the 10th (447 yards) at Miracle Hills Golf Club at Omaha, Nebraska, by Robert Mitera on October 7, 1965.

What is upwind distance?

UPWIND(aviation)-In airport traffic patterns leg paralel to the landing runway in the direction opposite the wind

Trending Questions
What is the probability that their next child will be male? What is the number of outcomes when one coin is flipped and eight sided die is rolled? What does per thousand population mean? There are 9 red and 6 blue marbles in a jar If you draw red marbles from the jar 7 times what is the experimental probability of drawing a red marble? When a continuous probability distribution is used to approximate a discrete probability distribution a value of 5 is added or subtracted from the area? If On a true and false test what is the probability of answering a question correctly if you make a random guess? How are odds related to probability? What is the function of sartorius? Two coins and one six sided cube are tossed together What is the probability of getting 2 heads and a four? Ninety-five percent of the parts coming off an assembly line are non-defective what is the probability that of 500 parts fewer than 472 are non-defective? Meaning on the tails side of a penny? Is 1.001 a probability? Which steps occurs last in the initiation phase of translation? A b c and d tells the truth with probability 13 suppose that a makes a statement and then d says that c says that b says that a was telling the truth what is the probability that a was true? Two six sided dice are rolled what is the probabiliy for the sum to be greater than 4? What is 4 5 as a percent? What are the odds of recovery from crack? Advantages and disadvantages of sampling techniques? What is the probability of Rr x Rr producing wrinkled seeds? When two dice are thrown simultaneously what is the probability that the sum of two numbers that turn up is less than 11?